find the domain of:
(A) f(x)= the square root of (-x squared - 4x + 5)
(B) f(x)= ln(4t-3)
(C) f(x)= 1 over (x to the third + 3x squared - x - 3)

Answer by stanbon(75887)

find the domain of:
(A) f(x)= sqrt(-x^2 - 4x + 5)
Solve: -x^2-4x+5 >= 0 for "x":
x^2 + 4x -5 <= 0
(x+5)(x-1) <= 0
Domain: -5 --------------------------
(B) f(x)= ln(4t-3)
Solve: 4t-3 > 0
4t > 3
Domain: t > 3/4
--------------------------
(C) f(x)= 1 over (x^3 + 3x^2 - x - 3)
I think your posting arrived garbled.
=============================
Cheers,
Stan H.